A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that solution A is 25% salt and Solution B is 50% salt. She wants to obtain 120 ounces of a mixture that is 35% salt. How many ounces of each solution should she use?
Answer by mananth(16946)

Percent ---------------- quantity
Solution A 25.00% ---------------- x ounces
SolutionB 50.00% ------ 120 - x ounces
Mixture 35.00% ---------------- 120
Total 120 ounces
25.00% x + 50.00% ( 120 - x ) = 35.00% * 120
25 x + 50 ( 120 - x ) = 4200
25 x + 6000 - 50 x = 4200
25 x - 50 x = 4200 - -6000
-25 x = -1800
/ -25
x = 72 ounces 25.00% Solution A
48 ounces 50.00% SolutionB

m.ananth@hotmail.ca